# U-shaped and Inverted-U Scaling behind Emergent Abilities of Large Language Models

### v1

@@ 4.1 Problem Formulation
We aim to predict the performance soar of traditional metrics before it happens. Specifically, we want to use only data before the emergent threshold to forecast the incidence of emergent abilities and the scaling trend past the emergent threshold. We compare the performance of our pipeline with the current iconic baseline of Sigmoid-based task-specific scaling law (Ye et al., 2023), which uses the Sigmoid function to regress accuracy on effective model size $M$.
@@ 4.2 Pipeline Overview
Figure 5: The overall pipeline of Slice-and-Sandwich. We slice questions into groups based on their difficulties, fit each group’s scaling trend, sandwich the overall performance to construct the scaling law on the linear metric, and finally project the scaling law back to the traditional metric.
Fig. 5 shows the overall pipeline of Slice-and-Sandwich. We use models smaller than the emergence threshold $T$ as the training set. As performance no longer stagnates with scale once we group questions by difficulty level, we fit the scaling trend of a continuous metric (binary Brier score in this paper) on the easiest question group and hardest question group separately and use the fitted scaling trend to forecast performance (measured in binary Brier Score) on easy and hard questions past $T$. We also use the training set to regress accuracy on the binary Brier Score and then use this estimated relation to convert the predicted binary Brier Score into predicted accuracy for models past the $T$.
@@ 4.3.1 Question Grouping
(a) MMLU.
(b) Arithmetic.
(c) Persian-QA.
Figure 6: U-shaped and inverted-U scaling on the MMLU, arithmetic, and Persian-QA datasets with group number $G=3$. U-shaped and inverted-U scaling trends still show.
To reduce data noise, we group questions into $G=3$ difficulty levels for Slice-and-Sandwich. Fig. 6 shows the binary Brier Score scaling trend for the easy, medium, and hard question groups. Notably, the scaling trend of the medium group resembles that of aggregate performance: it stagnates for a wide range of model sizes and then starts to improve sharply. By looking at the scaling trend for each difficulty level separately, the medium group’s pattern may simply be aggregating the scaling trend between easier and harder question groups.
@@
We use simple polynomial functions to fit the scaling trend of the binary Brier Score of the easy and hard question groups separately, using models before the emergence threshold $T$. We denote by $F^{c}_{e}(x)$ and $F^{c}_{h}(x)$ the fitted scaling trend of the easy and hard question groups, respectively, where $x$ is the effective model size. We then use $F^{c}_{e}(x)$ and $F^{c}_{h}(x)$ to forecast performance (measured in binary Brier Score) on the easy and the hard question groups of models with size $x$ above $T$.
We use the average of performance on the easy group and the hard group to forecast aggregated performance (measured in binary Brier Score):
$F^{c}(x)=\frac{1}{2}(F^{c}_{s}(x)+F^{c}_{h}(x))$ (6)
@@ 4.3.3 Obtaining Scaling Trend in Traditional Metric
Since what people usually care about ultimately are those traditional metrics such as accuracy (Hu et al., 2023), our last step is to project the forecast scaling trend in binary Brier Score, $F^{c}(x)$, back to scaling trend in accuracy, denoted by $F^{t}(x)$. One can replace binary Brier Score with other continuous metrics and accuracy with other traditional metrics. Specifically, we first estimate the relation between the continuous metric (binary Brier Score) and the traditional metric (accuracy) using models with sizes smaller than the emergent threshold $T$ as the training set. We denote the estimated mapping from binary Brier Score to accuracy as $G(\cdot)$.
Our forecast of scaling trend of accuracy is given by:
@@ 5.1 Fitting Scaling Trend of Easy Group and Hard Group
Figure 7: Data and polynomial fit for the easy and hard question groups on the MMLU, arithmetic, and Persian-QA datasets. The fitted trends encapsulate the actual trends.
@@
Figure 8: The binary-Brier-Score-based scaling law on the MMLU, arithmetic, and Persian-QA datasets acquired by taking the average of fitted trends of easy and hard question groups in Fig. 7.
Fig. 7 shows the fitted scaling trend of the easy and hard question groups on the MMLU, arithmetic, and Persian-QA datasets. Though the fitted scaling trend might either be underestimated, e.g., hard group of MMLU (Fig. 7(a)) and Persian-QA (Fig. 7(c)), or overestimated, e.g., easy group of MMLU (Fig. 7(a)), the actual scaling trend, they still encapsulate the overall trend. As a result, as shown in Fig. 8, averaging the scaling trend of the easy group and the hard group leads to pretty good prediction of the aggregate scaling trend.
5.2 Relation Between Accuracy and Binary Brier Score
@@
Figure 9: The relation between accuracy and the binary Brier Score on the MMLU, arithmetic, and Persian-QA datasets. The mapping function $G(x)$ from the binary Brier Score to accuracy can be well-modeled using small models as the training set.
Fig. 9 shows the close and almost linear relation between accuracy and binary Brier Score. As a result, simple ordinary least squares (OLS) regression using only models before the emergence threshold yields a precise mapping $G(\cdot)$ from binary Brier Score to accuracy.
@@ 5.3 Forecasting Scaling Trend in Accuracy
Figure 10: The accuracy-based scaling law on the MMLU, arithmetic, and Persian-QA datasets acquired by projecting the binary-Brier-Score-based scaling law back to accuracy-based scaling law by $G(x)$. Baseline is the Sigmoid-based regression (Owen, 2024).
Finally, Fig. 10 shows the accuracy-based scaling trend, $F^{t}(x)$, obtained through Eq. 7 with $G(\cdot)$, together with the baseline of fitting performance measured in accuracy on the Sigmoid function (Owen, 2024; Ruan et al., 2024). Instead of directly assuming a monotone scaling trend with the Sigmoid functional form, our Slice-and-Sandwich approach allows data to speak for itself. For the MMLU benchmark, our approach captures the forthcoming soaring trend, whereas the baseline approach does not. For the arithmetic dataset, though the baseline provides a seemingly decent forecast, it does not capture the sharp increase in the improvement speed at all, whereas our approach does. In short, our Slice-and-Sandwich approach could be more explainable and capable of capturing the soaring trends of emergent abilities. More experimental results on parameter robustness and the alternative method of Slice-and-Sandwich are in App. F.
6 Conclusions, Limitations, and Future Work
Conclusion. We propose to separately analyze scaling trends of LLM’s downstream performance for question groups with different difficulty levels. For six classical multiple-choice tasks with emergent abilities (three in the main paper and three in App. C), we demonstrate U-shaped scaling for hard questions and inverted-U scaling followed by steady improvement for easy questions. We propose the Slice-and-Sandwich pipeline to predict when emergent abilities will happen and forecast scaling trends afterward.
Limitations and Future Work. This work focuses on multiple-choice tasks. To apply our framework to other types of downstream tasks with emergent abilities, it is important and maybe not straightforward to identify a continuous metric that (1) differentiates easy questions from hard questions, and (2) is highly correlated with the traditional metric that people are ultimately interested in. We demonstrate this point in our preliminary analysis for string match tasks (App. G). We believe this to be a valuable avenue for future work.
@@ A.1 LLM Evaluation
We evaluate all datasets in this paper on the LM Evaluation Harness (Gao et al., 2024) platform. We adopt 56 models, including Gemma (Team et al., 2024), Llama (Touvron et al., 2023a), Llama-2 (Touvron et al., 2023b), RedPajama-INCITE (Computer, 2023), Yi (Young et al., 2024), StableLM (Stability-AI, 2023), MPT (Team, 2023), Falcon (Almazrouei et al., 2023), Pythia (Biderman et al., 2023), AMBER (Liu et al., 2023), Qwen (Bai et al., 2023), Qwen-1.5 (Bai et al., 2023), BLOOM (Workshop et al., 2022), DeepSeekMoE (Dai et al., 2024), OPT (Zhang et al., 2022), GPT-Neo (Black et al., 2021), Codegen (Nijkamp et al., 2023), XGLM (Lin et al., 2022b), and OpenLLaMA (Geng & Liu, 2023) families under FP16 precision. The evaluation time of each task varies from several hours to several days on 2 NVIDIA RTX A6000, depending on the question numbers and formats. We obtain each model’s effective model size (normalized training compute) through the released data by Ruan et al. (2024). We use $T=1.5,1.8$, and $2.3$ as the emergence threshold for the MMLU, arithmetic, and Persian-QA dataset, respectively. We calculate question difficulty level $q_{d}$ using models smaller than these thresholds. We adopt a 5-shot inference on the MMLU benchmark and a 2-shot inference on the arithmetic and Persian-QA datasets.
@@ A.2 Slice-and-Sandwich
We examine Slice-and-Sandwich on MMLU, arithmetic, and Persian-QA datasets with group number $G=3$. Models smaller than $T=1.5,1.8$, and $2.3$ in the MMLU, arithmetic, and Persian-QA datasets are the training set; other larger models are the testing set. We adopt polynomial regression to fit easy and hard question groups. Specifically, we adopt the polynomial order=5 and 2 for the easy and hard question groups, respectively.
Appendix B Discussion on Brier Score Definition
In the main paper, we use the model’s predicted probability of the correct choice conditional on available choices (see (4) to calculate the binary Brier Score. This section discusses the effect of such conditionalization. This section refers to the conditionalized binary Brier Score as the binary Brier Score we use in the main paper and the un-conditionalized binary Brier Score as the one without re-distributing output probabilities to available choices.
@@ Appendix B Discussion on Brier Score Definition
Figure A11: The un-conditionalized binary Brier Score vs. effective model size (M) on the MMLU, arithmetic, and Persian-QA datasets.
@@
Figure A12: The relation between accuracy and un-conditionalized binary Brier Score on the MMLU, arithmetic, and Persian-QA datasets.
Table A2: Comparison of correlation coefficients between accuracy and conditionalized/un-conditionalized binary Brier Score on the MMLU, arithmetic, and Persian-QA datasets. “P.”, “S.”, and “K.” stands for Pearson, Spearman, and Kendall, respectively. The conditionalized binary Brier Score has a consistently stronger correlation with accuracy.
@@
un-conditionalized 0.96 0.87 0.73 0.93 0.93 0.79 0.60 0.52 0.37
conditionalized 0.99 0.91 0.79 1.00 0.97 0.88 0.88 0.68 0.52
Fig. A11 shows the relationship between un-conditionalized binary Brier Score and effective model size (M) on all three datasets. For the MMLU dataset, model performance still exhibits flat scaling before the emergence threshold and sharp improvement past the emergence threshold. For the arithmetic and Persian-QA datasets, the scaling trend does not show a sharp increase and is easier to forecast performance under the un-conditionalized binary Brier Score past the emergence threshold. This is consistent with the finding of (Schaeffer et al., 2024b) that unconditionalized binary Brier Score is more correlated with the training compute than conditionalized binary Brier Score. However, Fig. A12 shows that the un-conditionalized binary Brier Score is not as closely related to accuracy as the conditionalized binary Brier Score for the arithmetic and especially the Persian-QA dataset. Table A2 corroborates this assertion by showing the correlation coefficient between accuracy and the conditionalized/un-conditionalized binary Brier Score.
@@
Figure A13: The U-shaped and inverted-U scaling with questions grouped and performances measured by the un-conditionalized binary Brier Score on the MMLU, arithmetic, and Persian-QA daatasets.
Fig. A13 shows the scaling trend by difficulty level for the un-conditionalized binary Brier Score. For the arithmetic and Persian-QA datasets, we no longer see inverse scaling on any intervals of effective Model size $M$. In fact, for both the arithmetic and the Persian-QA datasets, performance hovers around $-1$ on the hardest group, corresponding to the near-zero predicted probability of the correct choice. For hard questions, the model’s predicted probability on all available choices is close to zero. Therefore, without conditionalizing on available choices, we cannot differentiate between an initial random guess and the distracted phase at larger model sizes where the model places a higher probability on an available incorrect choice relative to the correct choice, which yields the U-shaped scaling of conditionalized binary Brier Score as discussed in the main paper.
@@ Appendix C Scaling Trend by Question Difficulty Level for Other Emergent Tasks
(a) Accuracy vs. effective model size (M).
(b) Binary Brier Score vs. effective model size (M).
@@
Figure A14: The accuracy, binary Brier Score, U-Shaped and inverted-U scaling on the Hindu knowledge dataset in BIG-bench (Srivastava et al., 2023).
(a) Accuracy vs. effective model size (M).
(b) Binary Brier Score vs. effective model size (M).
@@
Figure A15: The accuracy, binary Brier Score, U-Shaped and inverted-U scaling on the conceptual combinations dataset in BIG-bench (Srivastava et al., 2023).
@@
(a) Accuracy vs. effective model size (M).
(b) Binary Brier Score vs. effective model size (M).
@@
Figure A17: The accuracy, binary Brier Score, U-Shaped and inverted-U scaling on the analogical similarity dataset in BIG-bench (Srivastava et al., 2023).

### v2

@@ 4.1 Problem Formulation
We aim to predict the performance soar of traditional metrics before it happens. Specifically, we only use models before the emergent threshold $T$ to forecast the incidence of emergent abilities and the scaling trend after $T$. We compare the performance of our pipeline with the current iconic baseline of the Sigmoid-based task-specific scaling law (Ye et al., 2023), which uses the Sigmoid function to regress accuracy w.r.t. log compute $M$.
@@ 4.2 Pipeline Overview
Figure 6: The overall pipeline of Slice-and-Sandwich. We group questions into different difficulty levels, fit each group’s scaling trend, sandwich the overall performance to construct the scaling law on the linear metric, and finally project the scaling law back to the traditional metric.
Fig. 6 shows the overall pipeline of Slice-and-Sandwich. We use models smaller than the emergence threshold $T$ as the training set. As performance no longer stagnates with scale once we group questions by difficulty level, we fit the scaling trend of a continuous metric (TC Brier score in this paper) on the easiest question group and hardest question group separately and use the fitted scaling trend to forecast performance (measured in TC Brier Score) on easy and hard questions past $T$. We also use the training set to regress accuracy on the TC Brier Score and then use this estimated relation to convert the predicted TC Brier Score into predicted accuracy for models past the $T$.
@@ 4.3.1 Question Grouping
To reduce data noise, we group questions into $G=3$ difficulty levels, as in Fig. 4, for Slice-and-Sandwich and denote the level 1, 2 and 3 as easy, medium, and hard question groups. The medium group’s pattern is close to aggregating the scaling trend between easier and harder groups.
@@
We use simple polynomial regression to fit the scaling trend of the TC Brier Score of the easy and hard question groups separately using models before $T$. We denote by $F^{c}_{e}(x)$ and $F^{c}_{h}(x)$ the fitted scaling trend of the easy and hard question groups, respectively, where $x$ is the log compute. We then use $F^{c}_{e}(x)$ and $F^{c}_{h}(x)$ to forecast performance (measured in TC Brier Score) on the easy and the hard question groups of models with log compute $x$ above $T$.
We use the average of performance on the easy group and the hard group to forecast aggregated performance measured in TC Brier Score:
$F^{c}(x)=\frac{1}{2}(F^{c}_{e}(x)+F^{c}_{h}(x)),$ (5)
@@ 4.3.3 Obtaining Scaling Trend in Traditional Metric
Since people usually care about intuitive traditional metrics such as accuracy (Hu et al., 2023), our last step is to project the fitted scaling trend in TC Brier Score, $F^{c}(x)$, back to scaling trend in accuracy, denoted by $F^{t}(x)$. One can replace TC Brier Score with other continuous metrics and accuracy with other traditional metrics. Specifically, we first estimate the relation between the continuous metric (TC Brier Score) and the traditional metric (accuracy) using models with log computes smaller than $T$ as the training set. We denote the estimated mapping from TC Brier Score to accuracy as $G(\cdot)$. Our forecast of scaling trend of accuracy is given by:
@@ 5.1 Fitting Scaling Trend of Easy Group and Hard Group
Figure 7: Data and polynomial fit for the easy and hard question groups. The fitted trends encapsulate the actual trends.
@@
Figure 8: The TC-Brier-Score-based scaling law acquired by taking the average of fitted trends of easy and hard question groups in Fig. 7.
We adopt polynomial degree=2 and 5 for hard and easy questions, respectively, in response to our observation of U-shaped vs. inverted-U scaling. This parameter selection is based on our prior belief of polynomial regression’s fitting powers to fit the deep double descent and U-shaped scaling. Experimental results on parameter robustness are in App. G.
Fig. 7 shows the fitted scaling trend of the easy and hard question groups on the MMLU, arithmetic, and Persian-QA datasets. Empirically, fitted trends on hard questions are either precise or underestimated, e.g., hard group of MMLU (Fig. 7(a)) and arithmetic (Fig. 7(b)), due to lower fitting power of degree 2; fitted trends on easy questions are precise or overestimated, e.g., easy group of MMLU (Fig. 7(a)), due to a more considerable fitting power of degree 5. However, they still encapsulate the overall trend. Therefore, as shown in Fig. 8, taking their average can decrease the deviation and still lead to a precise prediction of the actual scaling trend.
5.2 Relation Between Accuracy and Target-Conditioned (TC) Brier Score
@@
Figure 9: The relation between accuracy and the TC Brier Score. The mapping function $G(x)$ from the TC Brier Score to accuracy can be well-modeled using small models as the training set.
Fig. 9 shows the close and almost linear relation between accuracy and TC Brier Score. As a result, simple ordinary least squares (OLS) regression using only models before the emergence threshold yields a precise mapping $G(\cdot)$ from TC Brier Score to accuracy.
@@ 5.3 Forecasting Scaling Trend in Accuracy
Figure 10: The accuracy-based scaling law acquired by projecting TC-Brier-Score-based scaling law back to accuracy-based scaling law by $G(x)$. Baseline is Sigmoid-based regression (Owen, 2024).
Fig. 10 shows the accuracy-based scaling trend, $F^{t}(x)$, obtained by Eq. 6 with $G(\cdot)$, together with the baseline of fitting the accuracy on the Sigmoid function (Owen, 2024; Ruan et al., 2024). Compared with the baseline, Slice-and-Sandwich better predicts and estimates the soaring performance by baking in more priors of the observed U-shaped and inverted-U scaling. For the MMLU, our approach captures the forthcoming soaring trend, whereas the baseline approach does not. For the arithmetic dataset, though the baseline provides a seemingly decent forecast, it does not capture the acceleration of performance increase, whereas our approach does. In short, our Slice-and-Sandwich approach is more explainable and capable of capturing the soaring trends of emergent abilities. A simple alternative method of Slice-and-Sandwich and its experimental results are in App. G.
6 Conclusions and Limitations
This work analyzes the LLM task-specific scaling law by grouping questions according to difficulty level. For six multiple-choice tasks with emergent abilities, we demonstrate U-shaped scaling for hard questions and inverted-U scaling followed by steady improvement for easy questions. These findings provide insight into the potential causes of emergent abilities. We then introduce the Slice-and-Sandwich pipeline to predict the emergence threshold and scaling law thereafter. However, as emergent phenomena are common across LLM benchmarks, it might be hard to claim all of them show clear U-shaped vs. inverted-U scaling. Furthermore, our focus is on multiple-choice tasks; applying our method to string-matching tasks requires identifying a continuous metric that differentiates easy questions from hard questions and correlates with the interested traditional metric. We illustrate and discuss this in App. H, highlighting future research direction.
@@ A.1 LLM Evaluation
We evaluate all datasets in this paper on the LM Evaluation Harness (Gao et al., 2024) platform. We adopt 56 models, including Gemma (Team et al., 2024), Llama (Touvron et al., 2023a), Llama-2 (Touvron et al., 2023b), RedPajama-INCITE (Computer, 2023), Yi (Young et al., 2024), StableLM (Stability-AI, 2023), MPT (Team, 2023), Falcon (Almazrouei et al., 2023), Pythia (Biderman et al., 2023), AMBER (Liu et al., 2023), Qwen (Bai et al., 2023), Qwen-1.5 (Bai et al., 2023), BLOOM (Workshop et al., 2022), DeepSeekMoE (Dai et al., 2024), OPT (Zhang et al., 2022), GPT-Neo (Black et al., 2021), Codegen (Nijkamp et al., 2023), XGLM (Lin et al., 2022b), and OpenLLaMA (Geng & Liu, 2023) families under FP16 precision. The evaluation time of each task varies from several hours to several days on 2 NVIDIA RTX A6000, depending on the question numbers and formats. We obtain each model’s log compute through the released data by Ruan et al. (2024). We use $T=1.5,1.8$, and $2.3$ as the emergence threshold for the MMLU, arithmetic, and Persian-QA dataset, respectively. We calculate question difficulty level $q_{d}$ using models smaller than these thresholds. We adopt 5-shot inference on the MMLU benchmark, 1-shot inference on the ARC and HellaSwag dataset, and 2-shot inference on Persian-QA, arithmetic, Hindu knowledge, conceptual combinations, analogical similarity, and abstract narrative understanding datasets (ARC, HellaSwag, and abstract narrative understanding datasets are used in App. D).
@@ A.2 Slice-and-Sandwich
In the main paper, we examine Slice-and-Sandwich on MMLU, arithmetic, and Persian-QA datasets with group number $G=3$. Models smaller than $T=1.5,1.8$, and $2.3$ in the MMLU, arithmetic, and Persian-QA datasets are the training set; other larger models are the testing set. We adopt polynomial regression to fit easy and hard question groups. Specifically, we adopt the polynomial order=5 and 2 for the easy and hard question groups, respectively.
Appendix B More Discussions on Brier Score
In the main paper, we use the model’s predicted probability of the correct class conditional on all classes to calculate the TC Brier Score (see Eq. 4). This section discusses the effect of such conditionalization. This section refers to the un-conditionalized TC Brier Score as the one without re-distributing output probabilities to all classes.
@@ Appendix B More Discussions on Brier Score
Figure A11: The un-conditionalized TC Brier Score vs. log compute (M) on the MMLU, arithmetic, and Persian-QA datasets.
@@
Figure A12: The relation between accuracy and un-conditionalized TC Brier Score on the MMLU, arithmetic, and Persian-QA datasets.
Table A2: Comparison of correlation coefficients between accuracy and TC Brier Score with and without conditionality on the MMLU, arithmetic, and Persian-QA datasets. “P.”, “S.”, and “K.” stands for Pearson, Spearman, and Kendall, respectively. The TC Brier Score with conditionality, i.e., the one we adopt in the main paper, has a consistently stronger correlation with accuracy.
@@
un-conditionalized TC Brier Score 0.96 0.87 0.73 0.93 0.93 0.79 0.60 0.52 0.37
TC Brier Score 0.99 0.91 0.79 1.00 0.97 0.88 0.88 0.68 0.52
Fig. A11 shows the relationship between un-conditionalized TC Brier Score and log compute $M$ on all three datasets. For the MMLU dataset, model performance still exhibits flat scaling before the emergence threshold and sharp improvement past the emergence threshold. For the arithmetic and Persian-QA datasets, the scaling trend does not show a sharp increase and is easier to forecast performance under the un-conditionalized TC Brier Score past the emergence threshold. This is consistent with the finding of (Schaeffer et al., 2024b) that the un-conditionalized measure is more correlated with the training compute than conditionalized ones. However, Fig. A12 shows that the un-conditionalized TC Brier Score is not as closely related to accuracy as the normal TC Brier Score for the arithmetic and especially the Persian-QA dataset. Table A2 corroborates this assertion by showing the correlation coefficient between accuracy and the normal/un-conditionalized TC Brier Score.
@@
Figure A13: The U-shaped and inverted-U scaling with questions grouped and performances measured by the un-conditionalized TC Brier Score on the MMLU, arithmetic, and Persian-QA daatasets.
Fig. A13 shows the scaling trend by difficulty level for the un-conditionalized TC Brier Score. For the arithmetic and Persian-QA datasets, we no longer see inverse scaling on any intervals of log compute $M$. In fact, for both the arithmetic and the Persian-QA datasets, performance hovers around $-1$ on the hardest group, corresponding to the near-zero predicted probability of the correct class. For hard questions, the model’s predicted probability on all classes is close to zero. Therefore, without conditionalizing on all classes, we cannot differentiate between an initial random guess and the distracted phase at larger model log computes where the model places a higher probability on an available incorrect class relative to the correct class, which yields the U-shaped scaling of normal TC Brier Score as discussed in the main paper.
@@ Appendix C Scaling Trend by Question Difficulty Level for Other Emergent Tasks
(a) Accuracy vs. log compute (M).
(b) TC Brier Score vs. log compute (M).
@@
Figure A14: The accuracy, TC Brier Score, U-Shaped and inverted-U scaling on the Hindu knowledge dataset in BIG-bench (Srivastava et al., 2023).
(a) Accuracy vs. log compute (M).
(b) TC Brier Score vs. log compute (M).
@@
Figure A15: The accuracy, TC Brier Score, U-Shaped and inverted-U scaling on the conceptual combinations dataset in BIG-bench (Srivastava et al., 2023).
@@
(a) Accuracy vs. log compute (M).
(b) TC Brier Score vs. log compute (M).
@@
Figure A17: The accuracy, TC Brier Score, U-Shaped and inverted-U scaling on the analogical similarity dataset in BIG-bench (Srivastava et al., 2023).